OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

Parth (parthkohli):

-_-

OpenStudy (anonymous):

(–4x)^0 = 1, but –4x^0 = –4.

Parth (parthkohli):

Do you mean to say \((-4x)^0 =-1\)?

OpenStudy (anonymous):

anything raised to the power of zero is 1

OpenStudy (anonymous):

so (-4x)^0 = 1 and you have (-4)(x^0) = (-4)(1) = -4

Parth (parthkohli):

Your PEMDAS says that you must first raise \(x\) to the power zero.\[-4 x ^0\]\[= -4 \times 1\]\[= -4\]

Parth (parthkohli):

But PEMDAS also says that you have to consider the parentheses first.\[(-4x)^0\]\[= (-4)^0 \times x^0\]\[=1 \times 1\]\[= 1\]

OpenStudy (ryan123345):

the parenthesis is basically telling u that u are taking the whole equation -4x and raising it to the 0 power which makes it 1. for the -4x0 u are actually raising the x to the 0 power making it 1 then u multiply the -4 to the 1 making it -4 :) u just have to pay attention to the parenthesis.